OpenStudy (anonymous):

can you reduce 3x+2y / 2x+y ?

OpenStudy (anonymous):

\[\frac{ 3x+2y }{ 2x+y }\]

Nnesha (nnesha):

rationalize the denominator??

Nnesha (nnesha):

what's the topic name ??

OpenStudy (anonymous):

just simplify

OpenStudy (anonymous):

i think it's just x+y but i duno i never trust myself lol i always need a snd opinion

Nnesha (nnesha):

O_O trust on ursefl

Nnesha (nnesha):

its already simplified :

OpenStudy (anonymous):

k

OpenStudy (anonymous):

like if it was \[\frac{ 4x+2y }{ 2x+y } \it could be \right?\]

OpenStudy (anonymous):

\[\frac{ 4x+2y }{ 2x+y }\]

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

k cool

Nnesha (nnesha):

yep :)

OpenStudy (anonymous):

sweeeeeeeeeeeeeeeeeeeeeeeeet

OpenStudy (anonymous):

tyyyyyyyyyyyyyyyy again

OpenStudy (anonymous):